OpenStudy (anonymous):

How do i find the distance between these points: (1,0,4) and (0,8,0) ??

OpenStudy (amistre64):

subtract square add sqrt

OpenStudy (amistre64):

its the same steps as finding the difference between any points in any vector space

OpenStudy (amistre64):

(1,0,4) -(0,8,0) -------- 1, -8,4 ; square em: 1, 64, 16 add em up 65 + 16 = 81 and sqrt it; sqrt(81) = 9 \[d=\sqrt{(x-xo)^2+(y-yo)^2+(z-zo)^2}\]

OpenStudy (anonymous):

Oh dang. Pretty easy. Thank you. (:

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

\[\sqrt{(0-1)^2+(8-0)^2+(0-4)^2}\] \[\sqrt{(-1)^2+(8)^2+(-4)^2}\] \[\sqrt{1+64+16}\] \[\sqrt{81}\] 9